# Machine-Learned Potentials for Solvation Modeling

### v2

@@ III.1.3   End-to-End NN-MLPs with Graph-Based Message Passing (MP-GNN-MLPs)
which encode cosine and cosine-squared dependencies on triplet angles. Here, $\hat{\mathbf{r}}_{ij}$ is the normalized direction vector and :::: denotes the double-dot product.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(l,\mathrm{scal})}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{scal}}(\mathbf{s}_{j}^{(l)},\mathbf{W}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (87)
$\mathbf{S}_{i\mathcal{N}(i)}^{(l,\mathrm{vec})}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{vec}}(\mathbf{s}_{j}^{(l)},\hat{\mathbf{r}}_{ij},\mathbf{W}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (88)
@@
ENINet Mao et al. (2024) is an $O(3)$-equivariant MP-GNN that encodes many-body geometric correlations using a line graph representation to explicitly model directional triplet interactions. It addresses the issue of message cancellation in equivariant GNNs with symmetric neighborhoods by leveraging explicit three-body angular context.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{local}}(\mathbf{h}_{j}^{(t-1)},\mathbf{r}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (97)
@@
NequIP Batzner et al. (2022) is an $E(3)$-equivariant MP-GNN that models atomic environments using geometric tensor features—scalars, vectors, and higher-order irreducible representations (irreps) of the 3D rotation group $O(3)$. It achieves high accuracy and data efficiency by explicitly enforcing symmetry across all layers.
@@
MACE Batatia et al. (2022) is an $E(3)$-equivariant MP-GNN that extends the atomic cluster expansion (ACE) into a learnable neural network framework using body-ordered tensor interactions. It operates on tensor-valued atom features and enforces exact equivariance via spherical harmonics and Clebsch–Gordan products.
@@
MACE achieves high accuracy and efficiency in modeling systems with strong angular dependencies, such as solvated molecules, interfaces, and complex condensed-phase materials.
@@
[…]
@@
$\mathbf{S}_{ij}^{\mathrm{scal},(l)}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{scal}}(x_{ik}^{(l)},\hat{\mathbf{r}}_{ik},w_{ik}^{(l)})\right\}_{k\in\mathcal{N}(i)},$ (113)
$\mathbf{S}_{ij}^{\mathrm{eq},(l)}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{eq}}(\mathbf{V}_{ik}^{(l)},Y_{\ell}(\hat{\mathbf{r}}_{ik}),w_{ik}^{(l)})\right\}_{k\in\mathcal{N}(i)},$ (114)
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{inv}}(f_{j}^{(t)},w_{ij})\right\}_{j\in\mathcal{N}(i)},$ (119)
$\mathbf{S}_{i\mathcal{N}_{\chi}(i)}^{(l,t)}$ =$=\left\{f_{\mathrm{msg}}^{(l)}(Y^{(l)}(\hat{\mathbf{r}}_{ij}),w_{ij})\right\}_{j\in\mathcal{N}(i)\cup\mathcal{N}_{\chi}(i)},$ (120)
@@
ViSNet Wang et al. (2024b) is an $SO(3)$-equivariant MP-GNN that uses a vector–scalar interactive message-passing mechanism (ViS-MP) to jointly update invariant and equivariant features. It replaces Clebsch–Gordan tensor contractions with efficient geometric encodings based on spherical harmonics and incorporates a runtime geometry calculation (RGC) module for bonded angular terms.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ =$=\left\{f_{\mathrm{msg}}^{\mathrm{inv}}(f_{j}^{(t)},w_{ij})\right\}_{j\in\mathcal{N}(i)},$ (127)
$\mathbf{S}_{i\mathcal{N}_{\chi}(i)}^{(l,t)}$ =$=\left\{f_{\mathrm{msg}}^{(l)}(Y^{(l)}(\hat{\mathbf{r}}_{ij}),w_{ij})\right\}_{j\in\mathcal{N}(i)\cup\mathcal{N}_{\chi}(i)},$ (128)
@@ III.2   Kernel models
Kernel models are strictly data-driven and, although they scale poorly with training set size compared to neural networks, they offer high accuracy in the low-data regime and can enforce desirable physical symmetries (e.g., permutational, rotational, and translational invariance) by construction. Force predictions are obtained via analytic derivatives of the energy, yielding energy-conserving and fully consistent gradients. These features make kernel-based potentials particularly attractive for solvation modeling tasks involving vibrational spectroscopy, conformer ranking, microsolvated clusters, or accurate force-driven dynamics in small to medium-sized systems.
@@
All kernel-based models can be paired with kernel functions—such as Gaussian, Laplacian, linear, or polynomial—that define a similarity measure between molecular or atomic environments. The energy is then predicted as a weighted sum over training samples
@@
Two widely used kernel functions are the Laplacian (also known as exponential) and Gaussian kernels, defined as
@@ III.3   Linear models
Linear-MLPs are linear and polynomial descriptor-based MLPs that use fixed, symmetry-adapted basis functions to approximate the potential energy surface. These methods rely on linear regression over carefully constructed features—such as tensor contractions, bispectrum components, or orthonormal polynomial expansions—to capture atomic environments in a manner consistent with physical symmetries. Unlike kernel methods, these models offer scalable training and faster predictions while maintaining systematic improvability and interpretability.
@@
Table 3 provides a comparative overview of the major MLP architectures used in solvation modeling. The models are categorized by learning paradigm—neural network–based potentials (NN-MLPs), kernel-based potentials (Kernel-MLPs), and linear models (Linear-MLPs)—and are further distinguished by their descriptor types and capacity to model long-range interactions.
@@ III.5   Gradient-domain models
Early kernel-based models, such as those based on GPR Li, Kermode, and De Vita (2015) and KRR Rupp, Ramakrishnan, and Von Lilienfeld (2015), predicted forces directly (i.e., these are F-MLPs) but did not guarantee integrability—i.e., the existence of a corresponding energy function. This challenge was addressed by sGDML Chmiela et al. (2019), which enforces symmetry and conservation analytically to produce high-fidelity force fields suitable for vibrational and thermodynamic modeling.
@@
NewtonNet Haghighatlari et al. (2022) is an $O(3)$-equivariant message-passing neural network inspired by Newton’s equations of motion. It introduces physically motivated modules—force, displacement, and energy calculators—that build latent equivariant features throughout the network, resulting in interpretable many-body interactions and improved data efficiency.
@@ III.6   Force-only models
$(\mathbf{R}\mathbf{r},\mathbf{R}\mathbf{F})\mapsto\text{training pair},$ (170)
@@
To predict scalar-valued properties such as core-level ionization energies and NMR chemical shifts, the authors use a modified version of the Coulomb matrix as a local atomic descriptor. Specifically, they construct local Coulomb matrices centered on each atom, consisting of the atom’s own nuclear charge and the Coulomb interactions with its $k$ nearest neighbors. These descriptors are explicitly made invariant to translation, rotation, and permutation of atom indices—a key requirement for learning scalar observables that do not depend on molecular orientation or labeling.
@@
$\mathbf{x}_{j}=\frac{Z_{j}}{\|\mathbf{r}_{j}-\mathbf{r}_{i}\|^{3}}(\mathbf{r}_{j}-\mathbf{r}_{i}),\quad j\neq i,\quad r_{rij}\leq\tau,$ (172)
@@
$\mathbf{C}_{i}=\sum_{j\neq i}\mathbf{x}_{j}\mathbf{x}_{j}^{\top},$ (173)
@@
$\mathbf{M}_{i}=\begin{bmatrix}Z_{j_{1}}&x^{\prime}_{j_{1}}-x_{i}&y^{\prime}_{j_{1}}-y_{i}&z^{\prime}_{j_{1}}-z_{i}\\ Z_{j_{2}}&x^{\prime}_{j_{2}}-x_{i}&y^{\prime}_{j_{2}}-y_{i}&z^{\prime}_{j_{2}}-z_{i}\\ \vdots&\vdots&\vdots&\vdots\\ Z_{j_{m}}&x^{\prime}_{j_{m}}-x_{i}&y^{\prime}_{j_{m}}-y_{i}&z^{\prime}_{j_{m}}-z_{i}\end{bmatrix},$ (176)
where $(x^{\prime}_{j}-x_{i},y^{\prime}_{j}-y_{i},z^{\prime}_{j}-z_{i})$ are the relative coordinates of neighbor atom $j$ projected into the local PCA-derived frame. The rows are sorted by increasing distance from the central atom $i$, enforcing permutation invariance with respect to atom indexing. The descriptor $\mathbf{M}_{i}$ includes only atoms $j$ within the cutoff sphere $r_{ij}\leq\tau$, where $m$ is the number of such neighbors.
@@
This PCA-based procedure thus ensures that the learned model respects the physical symmetries of the system. Although developed prior to modern GNNs, this approach anticipates many of the principles underlying equivariant architectures, including rotation- and translation-invariant descriptors and frame-consistent force prediction.
@@ III.7   Symmetry and equivariance
More generally, a function $f$ is invariant if $f(Tx)=f(x)$ and equivariant if $f(Tx)=Tf(x)$, where $T$ denotes a spatial transformation (e.g., rotation, translation, or reflection) acting on the input $x$. For out-of-sample predictions, MLPs should respect these symmetries by predicting invariant energies and equivariant forces, ensuring physically consistent behavior under arbitrary spatial configurations. This is achieved not by explicitly identifying transformations, but by embedding symmetry constraints into the model architecture, so that the outputs transform correctly—by design—even when the specific transformation was not encountered during training.
@@
[…]
These expressions describe how physical quantities behave under symmetry transformations applied to the molecular configuration =$\mathcal{R}=\{{\bf r}_{i}\}$.
@@
    The force $\mathbf{F}_{i}$ is a local vector quantity acting on atom $i$ and depends on the atomic configuration $\mathcal{R}$. When the entire system is rotated by a matrix $\mathbf{R}$, the force on each atom must also rotate accordingly: the vector itself changes, but the physical relation between geometry and force remains consistent. This transformation property—preserving directional relationships under symmetry operations—is the defining feature of equivariance.
@@
    The dipole moment $\boldsymbol{\mu}$, defined as the weighted sum of atomic positions $\mathbf{r}_{i}$ with partial charges $q_{i}$, $\boldsymbol{\mu}=\sum_{i}q_{i}\mathbf{r}_{i}$: transforms as a vector under rotations: $\boldsymbol{\mu}\to\mathbf{R}\boldsymbol{\mu}$. In the continuum limit, it is expressed as an integral over the charge density: $\boldsymbol{\mu}=\int\rho(\mathbf{r})\mathbf{r}\,d^{3}r$. More broadly, any global vector constructed from position-dependent quantities—such as total angular momentum or net force—exhibits the same equivariant behavior under spatial transformations.
@@
This rule mirrors angular momentum coupling in quantum mechanics Sakurai and Napolitano (2020): the composition of a degree-$\ell_{1}$ and degree-$\ell_{2}$ feature yields a range of angular components $\ell_{\text{out}}$ from $|\ell_{1}-\ell_{2}|$ up to $\ell_{1}+\ell_{2}$. Parity is multiplicative, so two even-parity features ($p=+1$) yield an even-parity output, whereas combining even and odd gives odd (+11=1+1$\cdot$-1=-1+ 1 ⋅ - 1 = - 1).

### v3

@@ III.1.3  End-to-End NN-MLPs with Graph-Based Message Passing (MP-GNN-MLPs)
which encode cosine and cosine-squared dependencies on triplet angles. Here, $\hat{\mathbf{r}}_{ij}$ is the normalized direction vector and :: denotes the double-dot product.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(l,\mathrm{scal})}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{scal}}(\mathbf{s}_{j}^{(l)},\mathbf{W}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (87)
$\mathbf{S}_{i\mathcal{N}(i)}^{(l,\mathrm{vec})}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{vec}}(\mathbf{s}_{j}^{(l)},\hat{\mathbf{r}}_{ij},\mathbf{W}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (88)
@@
ENINet Mao et al. (2024) is an $O(3)$-equivariant MP-GNN that encodes many-body geometric correlations using a line graph representation to model directional triplet interactions explicitly. It addresses the issue of message cancellation in equivariant GNNs with symmetric neighborhoods by leveraging explicit three-body angular context.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{local}}(\mathbf{h}_{j}^{(t-1)},\mathbf{r}_{ij})\right\}_{j\in\mathcal{N}(i)},$ (97)
@@
NequIP Batzner et al. (2022) is an $E(3)$-equivariant MP-GNN that models atomic environments using geometric tensor features such as scalars, vectors, and higher-order irreducible representations (irreps) of the 3D rotation group $O(3)$. It achieves high accuracy and data efficiency by explicitly enforcing symmetry across all layers.
@@
MACE Batatia et al. (2022) is an $E(3)$-equivariant MP-GNN that extends the ACE into a learnable neural network framework using body-ordered tensor interactions. It operates on tensor-valued atom features and enforces exact equivariance via spherical harmonics and Clebsch–Gordan products.
@@
MACE achieves high accuracy and efficiency in modeling systems with strong angular dependencies, such as solvated molecules, interfaces, and complex condensed-phase materials. The graph ACE (grACE), which generalizes ACE by employing graph-based, cluster basis functions to capture many-body interactions beyond the local atomic decomposition, exhibits superior accuracy to MACE while retaining a close formal connection to quantum-mechanical Hamiltonian expansions. Bochkarev, Lysogorskiy, and Drautz (2024)
@@
$x_{ij}^{(0)}$ $\in$ $\mathbb{R}^{C_{0}}\quad\text{(scalar feature)},$ (111)
$\mathbf{V}_{ij}^{(0)}$ $\in$ $\bigoplus_{\ell=1}^{\ell_{\max}}\mathbb{R}^{C_{\ell}}\otimes\mathcal{D}^{\ell}\quad\text{(tensor features)},$ (112)
@@
$\mathbf{S}_{ij}^{\mathrm{scal},(l)}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{scal}}(x_{ik}^{(l)},\hat{\mathbf{r}}_{ik},w_{ik}^{(l)})\right\}_{k\in\mathcal{N}(i)},$ (113)
$\mathbf{S}_{ij}^{\mathrm{eq},(l)}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{eq}}(\mathbf{V}_{ik}^{(l)},Y_{\ell}(\hat{\mathbf{r}}_{ik}),w_{ik}^{(l)})\right\}_{k\in\mathcal{N}(i)},$ (114)
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{inv}}(f_{j}^{(t)},w_{ij})\right\}_{j\in\mathcal{N}(i)},$ (119)
$\mathbf{S}_{i\mathcal{N}_{\chi}(i)}^{(l,t)}$ $=\left\{f_{\mathrm{msg}}^{(l)}(Y^{(l)}(\hat{\mathbf{r}}_{ij}),w_{ij})\right\}_{j\in\mathcal{N}(i)\cup\mathcal{N}_{\chi}(i)},$ (120)
@@
ViSNet Wang et al. (2024b) is an $SO(3)$-equivariant MP-GNN that uses a vector–scalar interactive message-passing mechanism (ViS-MP) to update invariant and equivariant features jointly. It replaces Clebsch–Gordan tensor contractions with efficient geometric encodings based on spherical harmonics and incorporates a runtime geometry calculation (RGC) module for bonded angular terms.
@@
$\mathbf{S}_{i\mathcal{N}(i)}^{(t)}$ $=\left\{f_{\mathrm{msg}}^{\mathrm{inv}}(f_{j}^{(t)},w_{ij})\right\}_{j\in\mathcal{N}(i)},$ (127)
$\mathbf{S}_{i\mathcal{N}_{\chi}(i)}^{(l,t)}$ $=\left\{f_{\mathrm{msg}}^{(l)}(Y^{(l)}(\hat{\mathbf{r}}_{ij}),w_{ij})\right\}_{j\in\mathcal{N}(i)\cup\mathcal{N}_{\chi}(i)},$ (128)
@@ III.2  Kernel models
Kernel models are strictly data-driven. While they scale poorly with training set size compared to neural networks, they offer high accuracy in the low-data regime. Further, they can enforce desirable physical symmetries (e.g., permutational, rotational, and translational invariance) by construction. Force predictions are obtained via analytic derivatives of the energy, yielding energy-conserving and fully consistent gradients. These features make kernel-based potentials particularly attractive for solvation modeling tasks involving vibrational spectroscopy, conformer ranking, microsolvated clusters, or accurate force-driven dynamics in small to medium-sized systems.
@@
All kernel-based models can be paired with kernel functions, such as Gaussian, Laplacian, linear, or polynomial, that define a similarity measure between molecular or atomic environments. The energy is then predicted as a weighted sum over training samples
@@
Two widely used kernel functions are the Laplacian (also known as exponential) and the Gaussian kernels, defined as
@@ III.3  Linear models
Linear-MLPs are linear and polynomial descriptor-based MLPs that use fixed, symmetry-adapted basis functions to approximate the potential energy surface. These methods rely on linear regression over carefully constructed features, such as tensor contractions, bispectrum components, or orthonormal polynomial expansions, to capture atomic environments in a manner consistent with physical symmetries. Unlike kernel methods, these models offer scalable training and faster predictions while maintaining systematic improvability and interpretability.
@@
Table 3 provides a comparative overview of the major MLP architectures used in solvation modeling. The models are categorized by learning paradigm: neural network–based potentials (NN-MLPs), kernel-based potentials (Kernel-MLPs), and linear models (Linear-MLPs). They are further distinguished by their descriptor types and capacity to model long-range interactions.
@@ III.5  Gradient-domain models
Early kernel-based models, such as those based on GPR Li, Kermode, and De Vita (2015) and KRR Rupp, Ramakrishnan, and Von Lilienfeld (2015), predicted forces directly (i.e., these are F-MLPs) but did not guarantee integrability, i.e., the existence of a corresponding energy function. This challenge was addressed by sGDML Chmiela et al. (2019), which enforces symmetry and conservation analytically to produce high-fidelity force fields suitable for vibrational and thermodynamic modeling.
@@
NewtonNet Haghighatlari et al. (2022) is an $O(3)$-equivariant message-passing neural network inspired by Newton’s equations of motion. It introduces physically motivated modules, such as force, displacement, and energy calculators, that build latent equivariant features throughout the network, resulting in interpretable many-body interactions and improved data efficiency.
@@ III.6  Force-only models
[…]
@@
To predict scalar-valued properties such as core-level ionization energies and NMR chemical shifts, the authors use a modified version of the Coulomb matrix as a local atomic descriptor. Specifically, they construct local Coulomb matrices centered on each atom, consisting of the atom’s own nuclear charge and the Coulomb interactions with its $k$ nearest neighbors. These descriptors are explicitly made invariant to translation, rotation, and permutation of atom indices, which is a key requirement for learning scalar observables that do not depend on molecular orientation or labeling.
@@
$\mathbf{x}_{j}=\frac{Z_{j}}{\|\mathbf{r}_{j}-\mathbf{r}_{i}\|^{3}}(\mathbf{r}_{j}-\mathbf{r}_{i}),\quad j\neq i,\quad r_{ij}\leq\tau,$ (172)
@@
$\mathbf{T}_{i}=\sum_{j\neq i}\mathbf{x}_{j}\mathbf{x}_{j}^{\top},$ (173)
@@
$\mathbf{M}_{i}=\begin{bmatrix}Z_{j_{1}}&x^{\prime}_{j_{1}}-x^{\prime}_{i}&y^{\prime}_{j_{1}}-y^{\prime}_{i}&z^{\prime}_{j_{1}}-z^{\prime}_{i}\\ Z_{j_{2}}&x^{\prime}_{j_{2}}-x^{\prime}_{i}&y^{\prime}_{j_{2}}-y^{\prime}_{i}&z^{\prime}_{j_{2}}-z^{\prime}_{i}\\ \vdots&\vdots&\vdots&\vdots\\ Z_{j_{m}}&x^{\prime}_{j_{m}}-x^{\prime}_{i}&y^{\prime}_{j_{m}}-y^{\prime}_{i}&z^{\prime}_{j_{m}}-z^{\prime}_{i}\end{bmatrix}$, (176)
where $(x^{\prime}_{j}-x^{\prime}_{i},y^{\prime}_{j}-y^{\prime}_{i},z^{\prime}_{j}-z^{\prime}_{i})$ are the relative coordinates of neighbor atom $j$ projected into the local PCA-derived frame. The rows are sorted by increasing distance from the central atom $i$, enforcing permutation invariance with respect to atom indexing. The descriptor $\mathbf{M}_{i}$ includes only atoms $j$ within the cutoff sphere $r_{ij}\leq\tau$, where $m$ is the number of such neighbors.
@@
This PCA-based procedure thus ensures that the learned model respects the physical symmetries of the system. Although developed before modern GNNs, this approach anticipates many of the principles underlying equivariant architectures, including rotation- and translation-invariant descriptors and frame-consistent force prediction.
@@ III.7  Symmetry and equivariance
More generally, a function $f$ is invariant if $f(Tx)=f(x)$ and equivariant if $f(Tx)=Tf(x)$, where $T$ denotes a spatial transformation (e.g., rotation, translation, or reflection) acting on the input $x$. For out-of-sample predictions, MLPs should respect these symmetries by predicting invariant energies and equivariant forces, ensuring physically consistent behavior under arbitrary spatial configurations. This is achieved not by explicitly identifying transformations, but by embedding symmetry constraints into the model architecture, so that the outputs transform correctly by design, even when the specific transformation was not encountered during training.
@@
Figure 10: Mean absolute errors (MAEs) from 50 out-of-sample predictions (training: 950 samples; test: 50) for energies (left panels) and forces (right panels), evaluated on the MD17 (A, B) and revised MD17 (C–F) datasets using various MP-GNN-MLPs. Panels A and B: MD17 molecules M1–M8: aspirin, benzene, ethanol, malonaldehyde, naphthalene, salicylic acid, toluene, and uracil. Panels C–F: revised MD17 molecules M1–M10: aspirin, azobenzene, benzene, ethanol, malonaldehyde, naphthalene, paracetamol, salicylic acid, toluene, and uracil. Methods (Panels A–B): SchNet, DimeNet, PhysNet, GemNet-Q, GemNet-T, PaiNN (force-only and joint energy/force variants), ENINet, SpookyNet, NequIP ($\ell=3$), NewtonNet, SO3krates, sGDML, ViSNet. Methods (Panels C–D): NequIP ($\ell=0$33), MACE, Allegro, FCHL19, GAP (SOAP), sGDML, ACE, MGNN, ViSNet. Panels E–F: Performance of NequIP for $\ell=0,1,2,3$. Only the top-performing models are highlighted in bold for clarity. The original data used to create this figure and other details are provided in the Supplementary Material.
These expressions describe how physical quantities behave under symmetry transformations applied to the molecular configuration $\mathcal{R}=\{{\bf r}_{i}\}$.
@@
    The force $\mathbf{F}_{i}$ is a local vector quantity acting on atom $i$ and depends on the atomic configuration $\mathcal{R}$. When the entire system is rotated by a matrix $\mathbf{R}$, the force on each atom must also rotate accordingly: the vector itself changes, but the physical relation between geometry and force remains consistent. This transformation property, preserving directional relationships under symmetry operations, is the defining feature of equivariance.
@@
    The dipole moment $\boldsymbol{\mu}$, defined as the weighted sum of atomic positions $\mathbf{r}_{i}$ with partial charges $q_{i}$, $\boldsymbol{\mu}=\sum_{i}q_{i}\mathbf{r}_{i}$: transforms as a vector under rotations: $\boldsymbol{\mu}\to\mathbf{R}\boldsymbol{\mu}$. In the continuum limit, it is expressed as an integral over the charge density: $\boldsymbol{\mu}=\int\rho(\mathbf{r})\mathbf{r}\,d^{3}r$. More broadly, any global vector constructed from position-dependent quantities, such as total angular momentum or net force, exhibits the same equivariant behavior under spatial transformations.
@@
This rule mirrors angular momentum coupling in quantum mechanics Sakurai and Napolitano (2020): the composition of a degree-$\ell_{1}$ and degree-$\ell_{2}$ feature yields a range of angular components $\ell_{\text{out}}$ from $|\ell_{1}-\ell_{2}|$ up to $\ell_{1}+\ell_{2}$. Parity is multiplicative, so two even-parity features ($p=+1$) yield an even-parity output, whereas combining even and odd gives odd ($1\cdot-1=-1$).